Hi! I keep doing this problem but it will not work for me and I don't know what I'm doing wrong.
[5(-x-8)-20] - [5-3(x-5)+19]
Can you show me the steps too?? Please and thank you. I need this by tonight if that would be alright.

Found 2 solutions by bucky, scott8148:
Answer by bucky(2189)

Given:
.
[5(-x-8)-20] - [5-3(x-5)+19]
.
To expand and simplify this begin by working each bracketed sets. The first bracketed
set is:
.
[5(-x-8)-20]
.
The terms inside the parentheses cannot be combined. So within the brackets do the multiplication
first. That is multiply out 5(-x-8) to get -5x -40. Substitute this result into the brackets
and you get:
.
[-5x - 40 -20]
.
Then the -40 and -20 are like terms that combine to -60. So the brackets now contain:
.
[-5x - 60]
.
Remember this result. Now let's go to the second set of brackets and work on it.
.
[5-3(x-5)+19]
.
Again the parentheses contain terms that cannot be combined. So the next step is to do the
distributed multiplication of -3 times the terms in the parentheses to get:
.
-3(x - 5) = -3x + 15
.
Substitute that into the brackets to get:
.
[5 - 3x + 15 + 19]
.
Now combine the numbers -3x + 5 + 15 + 19 = -3x + (5 + 15 + 19) = -3x + 39
.
This is the term in our second set of brackets. Now combining the two sets of brackets with
the negative sign between them results in:
.
[-5x - 60] - [-3x + 39]
.
Since the first set of brackets is preceded by an implied + sign, we can just remove the
brackets:
.
-5x - 60 - [-3x + 39]
.
And since the second set of brackets is preceded by a minus sign when we remove the brackets
we need to change the signs of the terms inside the brackets. This leads to:
.
-5x - 60 + 3x - 39
.
Now we can combine the -5x and + 3x to get -2x and combine the -60 and -39 to get - 99.
What we are left with is:
.
-2x - 99
.
That's the answer to this problem. Hope the explanation gives you some insight as to how
to do the problem.
.

Answer by scott8148(6628)

start by distributing with the inner parentheses ... be careful of signs ... [-5x-40-20]-[5-3x+15+19]

collect like terms [-5x-60]-[39-3x] ... distribute the minus sign -5x-60-39+3x

collect like terms again -2x-99